# Monge-Ampère gravity, optimal transport theory and their link to the Galileons

### v1

@@ II Mathematical background : the Monge-Ampère equation
which clearly demonstrates that the Poisson equation can be considered as a weak field limit of the MA equation. On the contrary, when $\beta\varphi_{\rm MA}\gg\mathbf{x}^{2}/2$ which is in some way a strong field limit, equation (3) approach equation (2). Therefore equation (3) can be seen as some sort of intermediate between the Poisson equation and the usual MA equation (2).
We can show fairly easily that equation (3) is translationally and rotationally invariant. However, this equation is no longer linear in $\varphi_{\rm MA}$ and scale-invariant contrary to the Poisson equation. Still, a characteristic scale doesn’t appear explicitly, as for example it does in the Yukawa coupling potential, since in this form $\alpha$ has the dimension of the inverse of a density, and $\beta$ has the dimension of a squared time. The original MA equation (2) is also affine invariant (Caffarelli, 2002). However it is not formally true for the potential $\varphi_{\rm MA}$ in equation (3), where affine-invariance is still present but modulo the harmonic potential.
@@ IV.1 Resolving MAG for spherically-symmetric densities
can be integrated for any spherically symmetric source. However, since we can only have an integral expression for $\partial_{r}\Phi$ with equation (13), we shall integrate from 00 to $r$, provided that the density profile is regular enough at $r=0$. Note that this is not the same as the usual practice for the Poisson equation which is integrated from $r$ to $\infty$, with the boundary condition that $\partial_{r}\varphi_{\rm N}\rightarrow\infty$ at $r\rightarrow 0$. With $\beta\varphi_{\rm MA}=\Phi-r^{2}/2$ we have:
@@ V.4 MAG beyond galactic scales
Figure 2: Thin slices of $256^{3}$ simulations in boxes of 60 Mpc/h with Newtonian dynamics (top panel) and Monge-Ampère gravity (bottom panel). The initial conditions of both simulations are identical and given by $\Lambda$CDM Planck Collaboration and Ade (2016) (See (Lévy et al., 2024)). The clustering is less prominent in Monge-Ampère because of screening at short distances. This is expected; GR as a spin-2 model of gravity is attractive, and any modification of GR, e.g. with a scalar degree of freedom, would typically yield a repulsive force.
The screening of MAG in strong field, has also been studied in a companion paper Lévy et al. (2024) where we have studied MAG as an effective theory at cosmological scales, using novel N-body simulations based on optimal transport theory. Fig. 2 demonstrates the suppression of clustering in MAG at small scales again due to this screening effect (see Lévy et al. (2024) for a full account). Indeed, modifications to GR typically yield a repulsive force, which explains the reason for the large interest in such theories as they provide an alternative, to the cosmological constant of GR, explanation of the accelerated expansion of the Universe.
We wish to make a comment at this point. Our analytic results and the snapshots in Fig. 2, clearly show that there are fewer small halos in MAG than in $\Lambda$CDM Universe. In this work, we have explained this feature in terms of a screening force and have shown that MAG is related to alternative models of gravity, e.g. Galileons, which naturally encapsulate similar screening mechanisms. However, many different mechanisms for suppression of small-scale structures have been proposed in the literature, among which is the Bose-Einstein or fuzzy dark matter model Hui et al. (2017) is which dark matter behaves as a super-fluid and obeys the Schrodinger equation. The decomposition of the wave function yields the continuity and Euler (or Hamilton-Jacobi for gradient velocities) equations, the latter with a quantum pressure term. The quantum pressure term suppresses small scale clustering. It is well-known that Schrodinger problem is directly related to optimal transport and hence connected to MAG equation (3) Léonard (2014). We shall explore this interesting aspect of MAG and its link to alternative models of dark matter in the forthcoming works.
@@ VI.1 Monge-Ampère equation as the field equation of quartic Galileon
$\mathcal{L}=\frac{1}{5}\mathcal{E}(\partial\partial\varphi)\varphi-\rho\varphi\,.$ (27)
@@
${\cal L}_{5}=-\frac{1}{5}\epsilon^{\mu\nu\rho\sigma}\epsilon^{\alpha\beta\gamma\delta}\varphi_{,\mu}\varphi_{,\alpha}\varphi_{,\nu\beta}\varphi_{,\rho\gamma}\varphi_{,\sigma\delta}-\rho\varphi\,,$ (29)
@@ VI.2 Monge-Ampère gravity (MAG) as the field equation of the sum of all Galileons in Minkowski space
Next we write a relativistic Lagrangian, in flat space, corresponding to the Monge-Ampère gravity equation used in MAG (3). We can replace the Laplacian of the identity map plus the Hessian in (3) by
$\partial_{\mu}\partial_{\nu}\varphi=\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi\,,$ (32)
which corresponds to relativistic MAG field equation, with a signature $\eta_{\mu\nu}=(-,+,+,+)$ :
$-\text{det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)\,=\rho\,.$ (33)
The corresponding Lagrangian density is:
${\cal L}=-\frac{1}{5}\psi\,{\rm det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)-\rho\,.$ (34)
We note that the Lagrangian (34) assumes that when we change the convention for the Minkowski metric, i.e. $\eta_{\mu\nu}\rightarrow-\eta_{\mu\nu}$ then we can also redefine $\psi\rightarrow-\psi$ which means that the determinant will become $(-1)^{4}$ times the Lagrangian density and hence the determinant and the Lagrangian are invariant under change of sign convention. Now one should verify that this Lagrangian does indeed correspond to the MAG equation. A small variation in the Lagrangian density yields :
$\delta{\cal L}$ $=$ $-\frac{1}{5}\delta\psi\,{\rm det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)$ (35)
-- $\frac{1}{5}\psi{\rm det}\,(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)(\eta_{\alpha\beta}+\partial_{\alpha}\partial_{\beta}\psi)^{-1}\partial^{\alpha}\partial^{\beta}\delta\psi$
-- $\rho\delta\psi$
$=$ $-\left({\rm det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)+\rho\right)\delta\psi\,.$ (36)
which yields the relativistic Monge-Ampère gravity equation of motion. We emphasis that the Lagrangian density (34) is not the Lagrangian of quintic Galileon.
@@
$\begin{split}-\text{det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)&=-\eta_{00}\text{det}(\mathbb{I}+D^{2}\psi)\\ &=+\text{det}(\mathbb{I}+D^{2}\psi)\\ \end{split}\,,$ (37)
which is the Monge-Ampère gravity equation in 3 dimensions that we had started with (3). The relativistic field $\psi$ then corresponds to the classical field $\varphi_{\rm MA}$ of MAG. Replacing the correct constant that we introduced for MAG, we can write the relativistic Lagrangian for MAG with correct constants as $-\frac{1}{5}\psi\text{det}(\eta_{\mu\nu}+(4\pi G_{\rm MA}\bar{\rho})^{-1}\partial_{\mu}\partial_{\nu}\psi)-\psi\rho/\bar{\rho}$.
Taking inspiration from the procedure we used in the previous subsection to derive MA equation from quartic Galileon, we see that the relativistic Lagrangian,
$\mathcal{L}=\psi\sum\text{det}_{3}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\psi)-\rho\psi\,,$ (38)
also corresponds to MAG, provided that we redefine $\psi\rightarrow-\psi$ when we change the sign of the metric $\eta_{\mu\nu}\rightarrow-\eta_{\mu\nu}$. Indeed, because of the presence of the identity in the MAG equation (3)–which is absent in the MA equation (2)–which translates to the Minkowski metric in the relativistic case, the static equation is the same in all dimension, and reduces from one form to the other when there is an invariance along a given dimension.
These Lagrangians involve the Minkowski metric $\eta_{\mu\nu}$ and may not be suited for a curved space-time. Replacing $\eta_{\mu\nu}$ by $g_{\mu\nu}$ would yield a new interaction between matter and space-time and furthermore one needs replace $\partial_{\mu}$ by $\nabla_{\mu}$. We shall not further investigate here the covariantisation of the above Lagrangian (38). Instead we shall seek a Lagrangian for MAG in terms of the Galileons.
In fact if we replace the $\varphi$ in the second derivatives in (29) by
$\varphi_{,\mu\nu}=\eta_{\mu\nu}+\psi_{,\mu\nu}$ (39)
We obtain MAG (3) as the equation of motion corresponding to a sum of all Galileons $\mathcal{L}_{i}$ Deffayet et al. (2009a), in flat space-time:
$\begin{split}{\cal L}_{\rm MAG}&=-\frac{1}{5}\epsilon^{\mu\nu\rho\sigma}\epsilon^{\alpha\beta\gamma\delta}\psi_{,\mu}\psi_{,\alpha}\big{(}\\ &(\psi_{,\nu\beta}+\eta_{\nu\beta})(\psi_{,\rho\gamma}+\eta_{\rho\gamma})(\psi_{,\sigma\delta}+\eta_{\sigma\delta})\big{)}-\rho\varphi\\ &=-\left(4!{\cal L}_{1}+4!{\cal L}_{2}+\frac{4!}{2!}{\cal L}_{3}+\frac{4!}{3!}{\cal L}_{4}+{\cal L}_{5}\right)-\rho\varphi\,.\end{split}$ (40)
It remains to be seen if the above procedure can be generalised to a curved background.
@@ VII Conclusion
The original 3-dimensional MA equation (2) has relativistic 4-dimensional equivalent, that arises in theories of Galileons and corresponds precisely to the static limit of the field equation of the Lagrangian of quartic Galileon. We have also shown that the MAG equation (3) (where the potential is written in terms of a harmonic potential) is the static limit of the field equation of the sum of the Lagrangians of all Galileons in Minkowski space.
@@
Acknowledgement We thank James Binney, Luc Blanchet, Gilles Esposito-Farèse, Bruno Lévy, André Lukas, Mohammad-Hossein Namdar, Farnik Nikakhtar, Felix Otto, Cale Rankin, Jean-Philippe Uzan, and Cliff Will for discussions.
@@ IX Appendix on the optimal transport
IX.1 Optimal mass transportation and the Monge-Ampère equation: two legacies of Gaspard Monge

### v2

@@ II Mathematical background : the Monge-Ampère equation
which might demonstrate that, the Poisson equation can be formally considered as a weak field limit of the MA equation. On the contrary, when $\beta\,D^{2}\varphi_{\rm MA}\gg 1$, for the same harmonic potential, then equation (3) approach equation (2). Therefore the MAG equation (3) can be seen as some sort of intermediate between the Poisson equation and the usual MA equation (2).
We can show fairly easily that equation (3) is translationally and rotationally invariant. However, this equation is no longer linear in $\varphi_{\rm MA}$ and scale-invariant contrary to the Poisson equation. Still, a characteristic scale appear through the combination of the mass of the source and the background density (as we shall show in expression (18)) and doesn’t appear in the usual manner, as for example it does in the Yukawa coupling potential, since in this form $\alpha$ has the dimension of the inverse of a density, and $\beta$ has the dimension of a squared time. The original MA equation (2) is also unimodular affine invariant (Caffarelli, 2002). However it is not formally true for the potential $\varphi_{\rm MA}$ in equation (3), where the same affine-invariance is still present but modulo the harmonic potential.
@@ IV.1 Resolving MAG for spherically-symmetric densities
can be integrated for any spherically symmetric source. However, since we can only have an integral expression for $\partial_{r}\Phi$ with equation (13), we shall integrate from 00 to $r$, provided that the density profile is regular enough at $r=0$. Note that this is not the same as the usual practice for the Poisson equation which is integrated from $r$ to $\infty$, with the boundary condition that $\partial_{r}\varphi_{\rm N}\rightarrow 0$ at $r\rightarrow\infty$. With $\beta\varphi_{\rm MA}=\Phi-r^{2}/2$ we have:
@@ V.4 MAG beyond galactic scales
Figure 2: Thin slices of $256^{3}$ simulations in boxes of 60 Mpc/h with Newtonian dynamics (top panel) and MAG (bottom panel) with $\Lambda$CDM initial conditions Planck Collaboration and Ade (2016) (See the companion paper (Lévy et al., 2024) where numerical simulations of MAG is presented). The clustering is less prominent in Monge-Ampère because of screening at short distances. This is expected; GR as a spin-2 model of gravity is attractive, and any modification of GR, e.g. with a scalar degree of freedom, would typically yield a repulsive force.
The screening of MAG in strong field, has also been studied in a companion paper Lévy et al. (2024) where we have studied MAG as an effective theory at cosmological scales, using novel N-body simulations based on optimal transport theory. Fig. 2 demonstrates the suppression of clustering in MAG at small scales again due to this screening effect (see Lévy et al. (2024) for a full account). Indeed, modifications to GR typically yield a repulsive force, which explains the reason for the large interest in such theories as they provide an alternative, to the cosmological constant of GR, explanation of the accelerated expansion of the Universe. We wish to make a comment at this point. Our analytic results and the snapshots in Fig. 2, clearly show that there are fewer small halos in MAG than in $\Lambda$CDM Universe. In this work, we have explained this feature in terms of a screening force and have shown that MAG is related to alternative models of gravity, e.g. Galileons, which naturally encapsulate similar screening mechanisms. However, many different mechanisms for suppression of small-scale structures have been proposed in the literature, among which is the Bose-Einstein or fuzzy dark matter model Hui et al. (2017) is which dark matter behaves as a super-fluid and obeys the Schrodinger equation. The decomposition of the wave function yields the continuity and Euler (or Hamilton-Jacobi for gradient velocities) equations, the latter with a quantum pressure term. The quantum pressure term suppresses small scale clustering. It is well-known that Schrodinger problem is directly related to optimal transport and hence connected to MAG equation (3) Léonard (2014). We shall explore this interesting aspect of MAG and its link to alternative models of dark matter in the forthcoming works.
@@ VI.1 Monge-Ampère equation as the field equation of quartic Galileon
$\mathcal{L}=\mathcal{E}(\partial\partial\varphi)\varphi-\rho\varphi\,.$ (27)
@@
${\cal L}_{5}=-\epsilon^{\mu\nu\rho\sigma}\epsilon^{\alpha\beta\gamma\delta}\varphi_{,\mu}\varphi_{,\alpha}\varphi_{,\nu\beta}\varphi_{,\rho\gamma}\varphi_{,\sigma\delta}-\rho\varphi\,,$ (29)
@@ VI.2 Monge-Ampère gravity (MAG) as the field equation of the sum of all Galileons in Minkowski space
Next we write a relativistic Lagrangian, in flat space, corresponding to the Monge-Ampère gravity equation used in MAG (3). We can formally replace the Laplacian of the identity map plus the Hessian in (3) by
$\partial_{\mu}\partial_{\nu}\varphi\rightarrow\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\varphi\,,$ (32)
while keeping the other derivatives the same, which corresponds to relativistic MAG field equation, with a signature $\eta_{\mu\nu}=(-,+,+,+)$ :
$-\text{det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\varphi)\,=\rho\,.$ (33)
@@
$\begin{split}-\text{det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\varphi)&=-\eta_{00}\text{det}(\mathbb{I}+D^{2}\varphi)\\ &=+\text{det}(\mathbb{I}+D^{2}\varphi)\\ \end{split}\,,$ (34)
which is the Monge-Ampère gravity equation in 3 dimensions that we had started with (3). The relativistic field $\varphi$ then corresponds to the classical field $\varphi_{\rm MA}$ of MAG. We note that the equation (33) assumes that when we change the convention for the Minkowski metric, i.e. $\eta_{\mu\nu}\rightarrow-\eta_{\mu\nu}$ then we can also redefine $\varphi\rightarrow-\varphi$ and hence the determinant is invariant under the sign convention.
Next, we seek a Lagrangian for MAG in terms of the Galileons. Next, we show that a particular linear combination of all the Galileons yield a Lagrangian forming this field equation (33) through Euler-Lagrange transformation. This Lagrangian is, using the notations of Deffayet et al. (2009a):
${\mathcal{L}}_{\rm MAG}=-\sum_{n=1}^{5}\frac{{\mathcal{L}}_{n}}{n!}-\rho\varphi\,,$ (35)
where we have defined:
${\mathcal{L}}_{n+1}=-A_{(2n)}\varphi_{,\mu_{1}}\varphi_{,\nu_{1}}\varphi_{,\mu_{2}\nu_{2}}...\varphi_{,\mu_{n}\nu_{n}}$ (36)
$A_{(2n)}=-\frac{1}{(5-n)!}\epsilon^{\mu_{1},...,\mu_{n},\alpha_{n+1},...,\alpha_{5}}\epsilon^{\nu_{1},...\nu_{n}}_{\alpha_{n+1},...,\alpha_{5}}\,.$ (37)
Note that the $2n$ indices ${\mu_{1},\nu_{1},...,\mu_{n},\nu_{n}}$ of the tensor $A_{(2n)}$ are not explicitly written for convenience. The variation of the Lagrangians ${\mathcal{L}}_{n+1}$ gives field equation ${\mathcal{E}}_{n+1}=0$, where:
${\mathcal{E}}_{n+1}=(n+1)A_{(2n)}\varphi_{,\mu_{1}\nu_{1}}...\varphi_{,\mu_{n}\nu_{n}}\,.$ (38)
Using these definitions, a simple way to derive that the Lagrangian (35) corresponds to the field equation (33) is to develop the determinant in this field equation:
$\begin{split}\text{det}(\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\varphi)&=\frac{1}{4!}\epsilon^{\mu_{1}...\mu_{4}}\epsilon^{\nu_{1}...\nu_{4}}\prod_{i=1}^{4}(\eta_{\mu_{i}\nu_{i}}+\partial_{\mu_{i}}\partial_{\nu_{i}}\varphi)\\ &=\frac{1}{4!}\sum_{n=1}^{5}C^{n-1}_{4}\frac{(5-n)!}{n}{\mathcal{E}}_{n}\\ &=\sum_{n=1}^{5}\frac{1}{n!}{\mathcal{E}}_{n}\,,\end{split}$ (39)
which shows that the elegant Lagrangian (35) yields the correct equation of motion. The above procedure can be generalized to a curved background. It is already well-known how to write the Lagrangian in curved space-time for Galileons Deffayet et al. (2009a) which would generate extra interactions with the curvature, however we shall not go further here as the effect is probably negligible for the numerical studies of MAG Lévy et al. (2024) that we are aiming at explaining in this work.
@@ VII Conclusion
The original 3-dimensional MA equation (2) has relativistic 4-dimensional equivalent, that arises in theories of Galileons and corresponds precisely to the static limit of the field equation of the Lagrangian of quartic Galileon. We have also shown that the MAG equation (3) (where the potential is written in terms of a harmonic potential) is the static limit of the field equation of a particular sum of the Lagrangians of all Galileons in Minkowski space, which can be generalised to curved space.
@@
Acknowledgement Special thanks go to Gilles Esposito-Farèse for numerous advices and discussions on Galileons. We thank James Binney, Luc Blanchet, Bruno Lévy, André Lukas, Mohammad-Hossein Namdar, Farnik Nikakhtar, Felix Otto, Cale Rankin, Jean-Philippe Uzan, and Cliff Will for discussions.
@@ IX Appendix on the optimal transport
IX.1 Optimal mass transportation and the Monge-Ampère equation: the two legacies of Gaspard Monge